OpenStudy (anonymous):

Provide a rough sketch of g(x). Label or identify the key features on the graph. How do I do this?

OpenStudy (anonymous):

Depends on what g(x) is. Right now your question is sort of like asking me to draw Bill, but I have no idea who Bill is or what he looks like.

OpenStudy (anonymous):

Idk, that's all it says on the question.

OpenStudy (anonymous):

g(x) is just another way of saying y.

OpenStudy (anonymous):

@shamil98 @agent0smith

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (anonymous):

okay @Ashleyisakitty

OpenStudy (anonymous):

heres the question: Heinz has a list of possible functions. Pick one of the g(x) functions below, show how to find the zeros, and then describe to Heinz the other key features of g(x). g(x) = x3 – x2 – 4x + 4 g(x) = x3 + 2x2 – 9x – 18 g(x) = x3 – 3x2 – 4x + 12 g(x) = x3 + 2x2 – 25x – 50 g(x) = 2x3 + 14x2 – 2x – 14 Provide a rough sketch of g(x). Label or identify the key features on the graph.

OpenStudy (ashleyisakitty):

Sorry, but I am poor at arithmetic.

OpenStudy (anonymous):

this is algebra lol but i dont understand what Provide a rough sketch of g(x). Label or identify the key features on the graph.

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@agent0smith any idea? what Provide a rough sketch of g(x). Label or identify the key features on the graph mean?

OpenStudy (anonymous):

@kirbykirby @Data_LG2

OpenStudy (anonymous):

@MDoodler @Whitemonsterbunny17

OpenStudy (anonymous):

@jim_thompson5910 @ParthKohli

whitemonsterbunny17 (whitemonsterbunny17):

Hmm, I'm sorry, but I have no idea how to do this. Maybe @paki can help?? (:

OpenStudy (anonymous):

ok @paki

OpenStudy (anonymous):

@karatechopper @Koikkara

OpenStudy (paki):

hm, sorry for late response, as at this time when you posted the question, there was night here... :) take the first function and to find the zeros of it: g(x)=x^3-x^2-4x+4 set function equal to 0, and then factor it. x^2(x-1) - 4(x -1) = 0 (x^2 -4)(x-1) = 0 (x+2)(x-2)(x-1) = 0 so x = 2, -2, and 1 Zeros of this function are: -2, 2, and 1... so x intercepts: -2, 2, and 1 y intercept: 4... The other key features of polynomial functions include end behavior, y intercept, and (when graphed) the axis of symmetry, and the vertex... for plotting the graph, please have a look here at the link... just put the values and it will give you the required graph.... https://www.desmos.com/calculator